The no.of integral values of m for which x coordinate of point of intersection of lines 3x+4y=9 & y=mx+1 is akso an integer is:

Grade:12

1 Answers

Arun
5 years ago
 
Solve both eqns simply and take out (x,y) coordinates.
put y = mx +1 in 3x + 4y – 9 = 0
U would get,
3x + 4(mx + 1) – 9 = 0
(3 + 4m)x = 6 
x = > 6/(3 + 4m)
and y = 6m/(3+4m) + 1 = > (7m +3)/(3+4m)
Now for x to be integer,
the modulus of denominator value should not be greater equals than 6.
Therefore,
(3+4m) less than equal to 6 and greater than -6
on solving m would satify as
m = 0, -1 .
simlarly do this for y coordinate.
 
Regards
Arun